9.3 Let w be a negative real number, z a 6
th root of w.
(a) Show that z (k) = ρ^
1
6
-

cos (pi+2kpi/6 )+ isin ( pi+2kpi/6), , k = 0, 1, 2, 3, 4, 5 is a formula for the
6th roots of w.
1
Expert's answer
2020-05-11T09:06:29-0400

(a) Let "w=\rho(\cos\phi+i\sin\phi)" and "z=r(\cos\theta+i sin\theta)" then by De Moivre's formula we have that "z^n=r^n(\cos n\theta+i \sin n\theta)"

"z^n=w \rightsquigarrow r^n(\cos n\theta+i \sin n\theta)=\rho(\cos\phi+i\sin\phi)"

Hence, "r^n=\rho, \,\, n\theta=\phi+2\pi k"

"w\in\mathbb R_- \rightsquigarrow \phi=\pi"

For "n=6\colon r=\rho^{1/6} \bigg( \cos \frac{\pi+2\pi k}{6} +i\sin \frac{\pi+2\pi k}{6} \bigg)"

for each "k\in\{0,1,\dots,n-1=5\}"